# Accelerating Hamiltonian Monte Carlo for Bayesian Inference in Neural Networks and Neural Operators

### v1

@@ 1 Introduction
MCMC methods comprise a set of algorithms to draw samples from arbitrary and intractable probability distributions. MCMC approaches can converge closely to the true posterior, but are computationally demanding. The potential for the MCMC technique, specifically the Hamiltonian Monte Carlo (HMC) technique, in the context of BNNs was explored by Neal [7]. Their method performed very well for small BNN models but failed to scale for larger models or datasets. To overcome this computational difficulty, Welling and Teh [17] introduced Stochastic Gradient Langevin Dynamics that improved scaling for large models. Following this, multiple stochastic gradient-based MCMC methods have been proposed in the literature, see [8, 9, 18, 19, 20, 21, 22, 23]. Despite these efficiency improvements, most of these methods are biased due to the lack of the Metropolis-Hasting correction step and subsampling of the data [24].
@@
A few works in the literature have explored accelerating HMC methods for neural networks, and even fewer have focused on combining VI and MCMC techniques. Salimans et al. [39] proposed a Markov chain variational inference (MCVI) method to include MCMC steps in the VI framework. Their goal, much like this work, was to combine the two methods to obtain computationally efficient approximations of the posterior distribution. However, unlike the present work, their work focused on incorporating MCMC into the optimization of the evidence lower bound (ELBO) in VI. Following this, Wolf et al [40] explored the use of HMC in the MCVI method. They introduced the Metropolis-Hasting acceptance step in the MCVI algorithm, ensuring convergence to the true posterior.
@@
The remaining sections of the manuscript are organized as follows. The VI and HMC methods used for posterior inference in Bayesian networks are briefly introduced in section 2. The proposed hybrid approach that combines VI and HMC is described in section 3. The efficiency and accuracy of the proposed VI-HMC method are demonstrated in section 4 via experiments on Bayesian neural networks (section 4.1) and Bayesian operator networks (section 4.2). In sections 4.1.1 and 4.1.2 the VI-HMC approach is demonstrated on two Bayesian neural networks to learn sinusoidal functions. A Bayesian operator network learning the solution of the Burgers equation using the hybrid VI-HMC approach is presented in section 4.2.1, and a Bayesian operator network that maps from upstream conditions to wall-pressure data on a cone in hypersonic flow is demonstrated in section 4.2.2. The computational cost of VI, HMC, and the proposed VI-HMC approach is compared in section 4.2.3. Limitations of the VI-HMC method are discussed in section 5 followed by conclusions in section 6.
@@ 2 Bayesian Inference in Neural Networks
The goal of a data-driven ML model is to learn a function $\boldsymbol{f}:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{y}}$ or an operator $\mathcal{G}:\mathcal{U}\rightarrow\mathcal{V}$ given a set of training data $\mathbb{D}$. Here, $\mathcal{U}$ and $\mathcal{V}$ are Banach spaces of vector-valued functions. When learning a function $\boldsymbol{f}$, the training data consists of a set of input-output points, $\mathbb{D}=\{x_{i},y_{i}\}_{i=1}^{N_{d}}$, where $x\in\mathbb{R}^{d_{x}}$ is the vector of inputs and $y\in\mathbb{R}^{d_{y}}$ is the vector of outputs. Similarly, while learning an operator $\mathcal{G}$, the data consist of input-output pairs, $\mathbb{D}=\{u_{i},v_{i}\}_{i=1}^{N_{d}}$, where $u\in\mathcal{U}$ is the input function and $v\in\mathcal{V}$ is the output function. The data driven ML model then learns an approximation $\mathcal{F}_{\boldsymbol{\Theta}}$ parameterized by $\boldsymbol{\Theta}\subset\mathbb{R}^{N_{\Theta}}$ for the function or the operator described above.
@@ 2.2 Hamiltonian Monte Carlo (HMC)
     $\boldsymbol{p}\leftarrow\mathcal{N}(0,\boldsymbol{M})$triangleright Sample momentum from a Gaussian distribution
     $\boldsymbol{\Theta}^{*},\boldsymbol{p}^{*}\leftarrow T_{s}(\boldsymbol{\Theta},\boldsymbol{p})$triangleright Solve the Hamiltonian equations (Eq. (8))
@@
         $u\leftarrow U[0,1]$triangleright Draw a sample from the uniform distribution
@@ 3 Accelerated HMC using VI sensitivities
We hypothesize that not all parameters of the network contribute equally to quantifying uncertainties in the neural network predictions and, moreover, that the contribution of each parameter to the prediction uncertainty can be assessed through a sensitivity analysis. Prediction uncertainties can then be effectively quantified by estimating the true joint distribution of a modest number of influential neural network parameters using HMC. The proposed VI-HMC method is conducted in three steps. In the first step, the distribution of the parameters is approximated using VI. In the second step, the sensitivities of the network predictions to each neural network parameter around the mean of VI parameters are estimated. In the third step, the network parameters are classified into two subsets—a low sensitivity and a high sensitivity subset—based on their contribution to the output variance. The parameters in the high-sensitivity subset are sampled using HMC, while the parameters in the low-sensitivity subsets are fixed at their mean values obtained from VI. This drastically reduces the dimension of the parameter space from which the samples need to be drawn during HMC, thereby accelerating HMC. This process is described in greater detail in Algorithm 2.
@@ 3.1 Sensitivity to network parameters
In VI, a Gaussian distribution $\mathcal{N}(\mu_{i},\sigma_{i})$ is learned for each network parameter $\theta_{i}\in\boldsymbol{\Theta}$, where $i,\ldots,N_{\Theta}$. Let us begin by considering the first-order Taylor series expansion of $\mathcal{F}_{\Theta}(x)$ evaluated at the mean parameter vector $\boldsymbol{\mu}$ given by
@@
Evaluating the variance of this expression yields
@@
Computing the sensitivities $S_{i}^{2}$ from Eq. (17), we then rank the parameters according to their sensitivities. Ordering the parameters from most sensitive to least sensitivity, we then truncate the summation to include $\hat{N}$ parameters based on an established threshold proportion of the variance captured, $\tau$, such that
$\hat{N}=\max\left\{N_{T}:\dfrac{\sum_{i=1}^{N_{T}}S_{i}^{2}}{\sum_{j=1}^{N_{\Theta}}S_{j}^{2}}\leq\tau\right\}$ (18)
@@ 3.2 HMC in reduced space
Samples from the distribution of the sensitive parameters $\mathscr{P}(\boldsymbol{\Theta}_{s}|\mathbb{D},\boldsymbol{\Theta}_{\sim s})$ are then drawn using HMC as shown in Algorithm 2.
@@
Compute: $\mathscr{P}(\boldsymbol{\Theta}|\mathbb{D})$triangleright Use VI to obtain approximate distributions of $\boldsymbol{\Theta}$
Require: Variance threshold, $\tau$triangleright Proportion of total variance to be included, Eq. (18)
Compute: $S_{i}^{2},\hat{S}$triangleright Compute the sensitivities and the sensitivity threshold
@@
Fix: $\boldsymbol{\Theta}_{\sim s}\leftarrow\mu_{\boldsymbol{\Theta}_{\sim s}}$triangleright Fix the parameters in $\boldsymbol{\Theta}_{\sim s}$ to their means
@@
     $\boldsymbol{p}\leftarrow\mathcal{N}(0,\boldsymbol{M})$triangleright Sample momentum from a Gaussian distribution
     $\boldsymbol{\Theta}^{*}_{s},\boldsymbol{p}^{*}\leftarrow T_{s}(\boldsymbol{\Theta}_{s},\boldsymbol{p})$triangleright Solve the Hamiltonian equations (Eq. (8))
@@
         $u\leftarrow U[0,1]$triangleright Draw a sample from the uniform distribution
@@ 3.3 Computational Benefits
The computational cost associated with drawing an independent sample from a $D$ dimensional distribution is approximately $\mathcal{O}(D^{5/4})$, considering the level of allowable error accumulation in the numerical integration of the Hamiltonian equations to achieve a reasonable acceptance rate [7, 49]. Therefore, reducing the dimension of the neural network’s parameter space will result in significant computational savings, which is the goal of the proposed hybrid VI-HMC approach. For example, reducing the parameter dimension from $10^{4}$ to $10^{3}$ corresponds to a nearly 20x reduction in expected computational cost. In other words, thanks to the reduced dimension, the Hamiltonian equations can be solved with a larger step size in the proposed VI-HMC approach, resulting in fewer time integration steps to convergence.
@@ 4 Numerical Examples
This section presents four numerical experiments that demonstrate the advantages of the proposed VI-HMC method. The first and second use BNNs to approximate functions from sparse data. The third and fourth examples apply Bayesian neural operators to learn complex partial differential equations. All numerical experiments were implemented in PyTorch [50]. The HMC algorithm uses the hamiltorch[25] package, and the VI-HMC method is implemented building on it.
@@ 4.1.1 Case I:
In the first case, the training data are generated using the parameters shown in Table 1 along with the noise parameter $\sigma_{d}=1E-3$. Twenty points with $x\in[-1,-0.2]\cup[0.2,1]$ are generated for training (with an intentional gap in the range $(-0.2,0.2)$) and 300 points with $x\in[-1.2,1.2]$ are generated for validation. A single hidden-layer neural network with two neurons and a sine activation function is considered. With the bias added to the hidden layer, the network has six parameters to learn that mimic the dimension of the parameterization in Eq. (21).
In the first step, the posterior distribution of the parameters is learned using VI with normal variational distribution. The network is trained until the ELBO loss converges, as seen in Figure 1(a). Also, shown in Figure 1(b) is the mean square error (MSE) convergence during VI training. Subsequently, the neural network parameter sensitivities are estimated as described in Section 3.1. Figure 2 shows the proportion of variance in the output captured for an increasing number of neural network parameters. The first four parameters are sufficient for a sensitivity threshold of $\tau=0.9$. Thus, in the VI-HMC step, the sample space reduces from 6 parameters to 4. The four most sensitive parameters are sampled using HMC, and the other two are fixed to their mean values obtained from VI.
The results obtained from learning the parameters using VI and standard HMC are compared with the results from the proposed VI-HMC method in Table 1. We observe that the parameters sampled through VI-HMC closely match those obtained using the full HMC approach. Note that the phase angles sampled by HMC are $\phi_{1}\approx 2\pi$, which is the same as $\phi_{1}=0$, and $\phi_{2}\approx 2.5\pi$, which is the same as $\phi_{2}=0.5\pi$.Also, Note that VI-HMC uses HMC to infer the distributions of the parameters $\omega_{1}$, $\phi_{2}$, $a$ and $b$ and selects $\omega_{2}$, and $\phi_{1}$ as their mean values from VI because their distributions contribute very little to the variance in the overall prediction.
@@
BNN predictions and associated uncertainties obtained by VI, HMC, and VI-HMC are presented in Figure 3. Note that VI overestimates the uncertainties compared to HMC whereas the uncertainty estimates of VI-HMC closely resemble the benchmark uncertainties obtained from HMC. Note also that the VI uncertainties are unrealistic in that they show high uncertainty even in regions with abundant data.
@@
Figure 3: Case I: BNN predictions and their uncertainties from (a) VI, (b) HMC, and (c) VI-HMC.
samples, mean prediction, training data, true function.

### v2

@@ 1 Introduction
MCMC methods comprise a set of algorithms to draw samples from arbitrary and intractable probability distributions. MCMC approaches can converge closely to the true posterior, but are computationally demanding. The potential for the MCMC technique, specifically the Hamiltonian Monte Carlo (HMC) technique, in the context of BNNs was explored by Neal [7]. Their method performed very well for small BNN models but failed to scale for larger models or datasets. To overcome this computational difficulty, Welling and Teh [17] introduced Stochastic Gradient Langevin Dynamics that improved scaling for large models. Following this, multiple stochastic gradient-based MCMC methods have been proposed in the literature, see [8, 9, 18, 19, 20, 21, 22, 23]. Despite these efficiency improvements, most of these methods are biased due to the lack of the Metropolis-Hastings correction step and subsampling of the data [24].
@@
A few works in the literature have explored accelerating HMC methods for neural networks, and even fewer have focused on combining VI and MCMC techniques. Salimans et al. [39] proposed a Markov chain variational inference (MCVI) method to include MCMC steps in the VI framework. Their goal, much like this work, was to combine the two methods to obtain accurate and computationally efficient approximations of the posterior distribution. Specifically, MCMC methods are used in MCVI to obtain an unbiased estimate of the gradient of the evidence lower bound (ELBO), which is then utilized in the optimization framework of VI to learn the variational parameters. In doing so, the optimization is improved, providing accurate estimates of the variational parameters while still maintaining the low cost benefit of VI methods. Following this, Wolf et al [40] explored and improved the use of HMC in the MCVI method. They introduced the Metropolis-Hastings acceptance step in the MCVI algorithm, ensuring convergence to the true posterior. Although these two works combined VI and HMC methods with the same goal as ours, they are fundamentally different from our hybrid approach. Unlike our work that focuses on reducing the computational cost of the HMC method using the information obtained from VI posteriors, MCVI [39, 40] focused on improving the accuracy of VI optimization by incorporating MCMC to compute gradients of ELBO.
@@
In the broader context of dimensionality reduction in Bayesian inference, several studies have focused on active subspaces methods. Cui et al [44] proposed a dimensionality reduction technique for Bayesian inverse problems by identifying a likelihood-informed subspace, determined by the relative influence of the prior and the likelihood on the posterior distribution. Constantine et al. [45] accelerated MCMC methods by finding an active subspace defined by the gradient of the negative log-likelihood. Following this, Tripathy and Bilionis [46] proposed a gradient free approach in formulating the active subspaces. Izmailov et al. [47] constructed low-dimensional parameter subspaces for inference in Bayesian deep learning and demonstrated performance improvements on Bayesian neural networks using different inference methods, including VI and No U-turn samplers. Along these lines, Daxberger et al. [48] proposed sampling a subnetwork to preserve the model uncertainties of the full network, where the subnetwork was selected based on the Wasserstein distance between the approximate posterior for the full network and the approximate posterior for the subnetwork. While these efforts, along with other developments in active subspaces [49, 50, 51, 52], pursue the common objective of efficient Bayesian inference through dimensionality reduction, the methods they adopt differ from the approach proposed in this work.
The remaining sections of the manuscript are organized as follows. The VI and HMC methods used for posterior inference in Bayesian networks are briefly introduced in section 2. The proposed hybrid approach that combines VI and HMC is described in section 3. The efficiency and accuracy of the proposed VI-HMC method are demonstrated in section 4 via experiments on Bayesian neural networks (section 4.1) and Bayesian operator networks (section 4.2). In sections 4.1.1 and 4.1.2, the VI-HMC approach is demonstrated on two Bayesian neural networks to learn sinusoidal functions. A Bayesian operator network learning the solution of the Burgers equation using the hybrid VI-HMC approach is presented in section 4.2.1, and a Bayesian operator network that maps from upstream conditions to wall-pressure data on a cone in hypersonic flow is demonstrated in section 4.2.2. The computational cost of VI, HMC, and the proposed VI-HMC approach is compared in section 4.2.3. Limitations of the VI-HMC method are discussed in section 5, followed by conclusions in section 6.
@@ 2 Bayesian Inference in Neural Networks
The goal of a data-driven ML model is to learn a function $\boldsymbol{f}:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{y}}$ or an operator $\mathcal{G}:\mathcal{U}\rightarrow\mathcal{V}$ given a set of training data $\mathbb{D}$. Here, $\mathcal{U}$ and $\mathcal{V}$ are Banach spaces of vector-valued functions. When learning a function $\boldsymbol{f}$, the training data consists of a set of input-output points, $\mathbb{D}=\{x_{i},y_{i}\}_{i=1}^{N_{d}}$, where $x\in\mathbb{R}^{d_{x}}$ is the vector of inputs and $y\in\mathbb{R}^{d_{y}}$ is the vector of outputs. Similarly, while learning an operator $\mathcal{G}$, the data consist of input-output pairs, $\mathbb{D}=\{u_{i},v_{i}\}_{i=1}^{N_{d}}$, where $u\in\mathcal{U}$ is the input function and $v\in\mathcal{V}$ is the output function. The data-driven ML model then learns an approximation $\mathcal{F}_{\boldsymbol{\Theta}}$ parameterized by $\boldsymbol{\Theta}\subset\mathbb{R}^{N_{\Theta}}$ for the function or the operator described above.
@@ 2.2 Hamiltonian Monte Carlo (HMC)
  $\boldsymbol{p}\leftarrow\mathcal{N}(0,\boldsymbol{M})$ $\triangleright$ Sample momentum from a Gaussian distribution
  $\boldsymbol{\Theta}^{*},\boldsymbol{p}^{*}\leftarrow T_{s}(\boldsymbol{\Theta},\boldsymbol{p})$ $\triangleright$ Solve the Hamiltonian equations (Eq. (8))
@@
   $u\leftarrow U[0,1]$ $\triangleright$ Draw a sample from the uniform distribution
@@ 3 Accelerated HMC using VI sensitivities
We hypothesize that not all parameters of the network contribute equally to quantifying uncertainties in the neural network predictions and, moreover, that the contribution of each parameter to the prediction uncertainty can be assessed through a sensitivity analysis. Prediction uncertainties can then be effectively quantified by estimating the true joint distribution of a modest number of influential neural network parameters using HMC. The proposed VI-HMC method is conducted in three steps. In the first step, the distribution of the parameters is approximated using VI. In the second step, the sensitivities of the network predictions to each neural network parameter around the mean of the VI parameters are estimated. In the third step, the network parameters are classified into two subsets—a low sensitivity and a high sensitivity subset—based on their contribution to the output variance. The parameters in the high-sensitivity subset are sampled using HMC, while the parameters in the low-sensitivity subsets are fixed at their mean values obtained from VI. This drastically reduces the dimension of the parameter space from which the samples need to be drawn during HMC, thereby accelerating HMC. This process is described in greater detail in Algorithm 2.
@@ 3.1 Sensitivity to network parameters
In VI, a Gaussian distribution $\mathcal{N}(\mu_{i},\sigma_{i})$ is learned for each network parameter $\theta_{i}\in\boldsymbol{\Theta}$, where $i,\ldots,N_{\Theta}$, and $\mu_{i}$, $\sigma_{i}$ are the mean and the standard deviation of the parameter, respectively. Let us begin by considering the first-order Taylor series expansion of $\mathcal{F}_{\Theta}(x)$ evaluated at the mean parameter vector $\boldsymbol{\mu}$ given by
@@
Here, $\mathcal{F}_{\boldsymbol{\mu}}(x)$ is the neural network or the neural operator evaluated at the mean parameters $\mu$ for an input $x$. Evaluating the variance of this expression yields
@@
Computing the sensitivities $S_{i}^{2}$ from Eq. (17), we then rank the parameters according to their sensitivities. Ordering the parameters from most to least sensitive, we then truncate the summation to include $\hat{N}$ parameters based on an established threshold proportion of the variance captured, $\tau$, such that
$\hat{N}=\max\left\{N_{T}:\dfrac{\sum_{i=1}^{N_{T}}S_{i}^{2}}{\sum_{j=1}^{N_{\Theta}}S_{j}^{2}}\leq\tau\right\}.$ (18)
That is, the number of sensitive parameters $\hat{N}$ is the maximum number that needs to be included in the summation such that the variance ratio is less than or equal to $\tau$. Although a first-order Taylor series expansion is proposed in Eq.(13), higher order terms can be included, resulting in a modification to the sensitivity scores in Eq. (17). These changes can be incorporated into the current framework without modifying the rest of the hybrid VI-HMC approach.
@@ 3.2 HMC in reduced space
Samples from the distribution of the sensitive parameters $\mathscr{P}(\boldsymbol{\Theta}_{s}|\mathbb{D},\boldsymbol{\mu}_{\sim s})$ are then drawn using HMC as shown in Algorithm 2.
@@
Compute: $\mathscr{P}(\boldsymbol{\Theta}|\mathbb{D})$ $\triangleright$ Use VI to obtain approximate distributions of $\boldsymbol{\Theta}$
Require: Variance threshold, $\tau$ $\triangleright$ Proportion of total variance to be included, Eq. (18)
Compute: $S_{i}^{2},\hat{S} \triangleright$ Compute the sensitivities and the sensitivity threshold
@@
Fix: $\boldsymbol{\Theta}_{\sim s}\leftarrow\mu_{\boldsymbol{\Theta}_{\sim s}}$ $\triangleright$ Fix the parameters in $\boldsymbol{\Theta}_{\sim s}$ to their means
@@
  $\boldsymbol{p}\leftarrow\mathcal{N}(0,\boldsymbol{M})$ $\triangleright$ Sample momentum from a Gaussian distribution
  $\boldsymbol{\Theta}^{*}_{s},\boldsymbol{p}^{*}\leftarrow T_{s}(\boldsymbol{\Theta}_{s},\boldsymbol{p})$ $\triangleright$ Solve the Hamiltonian equations (Eq. (8))
@@
   $u\leftarrow U[0,1]$ $\triangleright$ Draw a sample from the uniform distribution
@@ 3.3 Computational Benefits
The computational cost associated with drawing an independent sample from a $D$-dimensional distribution is approximately $\mathcal{O}(D^{5/4})$, considering the level of allowable error accumulation in the numerical integration of the Hamiltonian equations to achieve a reasonable acceptance rate [7, 58]. Therefore, reducing the dimension of the neural network’s parameter space will result in significant computational savings, which is the goal of the proposed hybrid VI-HMC approach. For example, reducing the parameter dimension from $10^{4}$ to $10^{3}$ corresponds to a nearly 20x reduction in expected computational cost. In other words, thanks to the reduced dimensionality, the Hamiltonian equations can be solved with a larger step size in the proposed VI-HMC approach, resulting in fewer time integration steps to convergence.
@@ 4 Numerical Examples
This section presents four numerical experiments that demonstrate the advantages of the proposed VI-HMC method. The first and second use BNNs to approximate functions from sparse data. The third and fourth examples apply Bayesian neural operators to learn complex partial differential equations. All numerical experiments were implemented in PyTorch [59]. VI layers are implemented following [60]. The HMC algorithm uses the hamiltorch[25] package, and the VI-HMC method is implemented by building on it. All the hyper-parameters used in the experiments are summarized in C.
Code to implement the proposed VI-HMC approach can be found at this GitHub repository: https://github.com/ponkrshnan/VI-HMC.git
@@ 4.1.1 Case I:
In the first case, the training data are generated using the parameters shown in Table 1 along with the noise parameter $\sigma_{d}=1E-3$. Twenty points with $x\in[-1,-0.2]\cup[0.2,1]$ are generated for training (with an intentional gap in the range (0.2,0.2)(-0.2,0.2)) and 300 points with $x\in[-1.2,1.2]$ are generated for validation. A single hidden-layer neural network with two neurons and a sine activation function is considered. With the bias added to the hidden layer, the network has six parameters to learn that mimic the dimension of the parameterization in Eq. (21).
In the first step, the posterior distribution of the parameters is learned using VI with a normal variational distribution. The network is trained until the ELBO loss converges, as seen in Figure 1(a). Also, shown in Figure 1(b) is the mean square error (MSE) convergence during VI training. Subsequently, the neural network parameter sensitivities are estimated as described in Section 3.1. Figure 2 shows the proportion of variance in the output captured for an increasing number of neural network parameters. The first four parameters are sufficient for a sensitivity threshold of $\tau=0.9$. Thus, in the VI-HMC step, the sample space reduces from 6 parameters to 4. The four most sensitive parameters are sampled using HMC, and the other two are fixed to their mean values obtained from VI.
The results obtained from learning the parameters using VI and standard HMC are compared with the results from the proposed VI-HMC method in Table 1. We observe that the parameters sampled through VI-HMC closely match those obtained using the full HMC approach. Note that the phase angles sampled by HMC are $\phi_{1}\approx 2\pi$, which is the same as $\phi_{1}=0$, and $\phi_{2}\approx 2.5\pi$, which is the same as $\phi_{2}=0.5\pi$.Also, note that VI-HMC uses HMC to infer the distributions of the parameters $\omega_{1}$, $\phi_{2}$, $a$, and $b$ and selects $\omega_{2}$ and $\phi_{1}$ as their mean values from VI because their distributions contribute very little to the variance in the overall prediction. The sensitivity rankings of the parameters are verified by computing analytical derivatives of the sinusoidal function in Eq. (21) with respect to the six parameters and computing the sensitivity scores according to Eq. (17).
@@
BNN predictions and associated uncertainties obtained by VI, HMC, and VI-HMC are presented in Figure 3. Note that VI overestimates the uncertainties compared to HMC, whereas the uncertainty estimates of VI-HMC closely resemble the benchmark uncertainties obtained from HMC. Note also that the VI uncertainties are unrealistic in that they show high uncertainty even in regions with abundant data.
@@
Figure 3: Case I: BNN predictions and their uncertainties from (a) VI, (b) HMC, and (c) VI-HMC. samples, mean prediction, training data, true function.